Senior Life Co. is an HMO for businesses in the Portland area. The following account balances appear on the balance sheet of Senior Life Co.: Common stock (360,000 shares authorized; 4,000 shares issued), $125 par, $500,000; Paid-In Capital in excess of par— common stock, $100,000; and Retained earnings, $6,000,000. The board of directors declared a 1% stock dividend when the market price of the stock was $153 a share. Senior Life Co. reported no income or loss for the current year.

If an amount box does not require an entry, leave it blank. If no entry is required, select "No entry required" from the dropdown.

a1. Journalize the entry to record the declaration of the dividend, capitalizing an amount equal to market value.

a2. Journalize the entry to record the issuance of the stock certificates.

b. Determine the following amounts before the stock dividend was declared: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $
Total retained earnings $
Total stockholders' equity $

c. Determine the following amounts after the stock dividend was declared and closing entries were recorded at the end of the year: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

Total paid-in capital $
Total retained earnings $
Total stockholders' equity $
Answer #1
a1
Stock Dividends 6120 =4000*1%*153
     Stock Dividends Distributable 5000 =4000*1%*125
     Paid in Capital in Excess of par-Common Stock 1120
a2
Stock Dividends Distributable 5000
     Common Stock 5000
b
Total paid-in capital 600000 =500000+100000
Total retained earnings 6000000
Total stockholders' equity 6600000
c
Total paid-in capital 606120 =600000+6120
Total retained earnings 5993880 =6000000-6120
Total stockholders' equity 6600000